OpenStudy (anonymous):

Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21.

OpenStudy (anonymous):

hi welcome to openstudy! just to make sure, it is 6(x+2) right? not 6(x+2)^2?

OpenStudy (pawanyadav):

Y=f(x) Y+3=6(x+2) (Y+3)/6=x+2 (y+3)/6-2=x Interchange x and y (x+3)/6-2=y Given x=21 as now above function is inverse. (21+3)/6-2=y 4-2=y y=2. answer